# Simulating and Understanding Deceptive Behaviors in Long-Horizon Interactions

### v2

@@ B.1 Experimental Details
All models were evaluated under a consistent configuration of the long-horizon simulation framework. The supervisor agent was instantiated with GPT-5-mini, chosen to balance reasoning ability and response efficiency, with its satisfaction threshold fixed at $\tau=0.2$ and memory $K=7$. The event triggering probability was set to $p=0.5$ in our main experiment. These parameters were chosen to balance task difficulty and environmental stress, ensuring that trajectories are neither trivially easy (leading to minimal opportunities for deception) nor prohibitively difficult (leading to uniform failure). We implement the deception auditor using GPT-5 (with reasoning_effort=high), due to its strong reasoning and evaluation capabilities. Importantly, GPT-5 (OpenAI, 2025a) itself is not included among the models under evaluation in the main experiment, in order to avoid potential bias arising from having the same model both generate and audit behavior.
@@
During the simulation, the event system is triggered to sample an event with probability $p_{\text{event}}$. Once the system is triggered, it uniformly samples event category, event name, and pressure level, then selects the corresponding event. The sampler is initialized with a fixed seed to guarantee reproducibility and fair cross-model comparison. Given a task index and its project phase, the sampler deterministically produces the same event $e$ and thus the event trajectory for a task stream is deterministic. Running different models under the same event seed, therefore, exposes them to an identical sequence of event realizations, guaranteeing identical experimental conditions.
@@ Appendix C Case Study
In this section, we show cases that were used to conduct qualitative analysis. Each case is provided alongside a paragraph of detailed analysis.
C.1 Deception in Long-horizon Tasks
C.1.1 Case 1. Chain of Deception
@@ C.1 Deception in Long-horizon Tasks
C.1.2 Case 2. Repetitive Deceptive Behaviors
@@
C.2 Behaviors Under Critical Pressure

### v3

@@ A.5 Deception Auditor
A.6 Limitations and Future Work
LH-Deception focuses on systematically quantifying emergent deception in long-horizon, language-based interactions. We acknowledge two primary limitations, which also serve as critical directions for future research.
1. Action Space: A key limitation of our current framework is that the performer’s action space is intentionally confined to natural language generation, designed to first isolate, measure, and understand the dynamics of social and strategic deception (e.g., falsification, concealment, and equivocation) in a controlled setting.
However, we acknowledge that many complex, real-world long-horizon tasks involve a much larger action space, particularly the use of tools (e.g., API calls, code execution, web searches). The integration of tools introduces new, complex vectors for deception that our current framework does not capture. For example, an agent could deceptively misreport a tool’s output, strategically conceal tool-related errors, or even tamper with its own API calls to produce a misleading result. Extending our simulation framework to include tool use is a crucial and exciting next step to capture these more sophisticated, technically-grounded deceptive strategies.
2. “Black-Box” Auditing: A second limitation pertains to our black-box auditing design. The Deception Auditor intentionally reviews only the observable history (i.e., the performer’s final, external outputs) rather than its internal chain-of-thought (CoT). This choice is a simulation of a realistic and critical accountability scenario, where an end-user, manager, or external regulator must judge a system’s trustworthiness based solely on its actions, without access to its internal thoughts.
However, this design means our framework cannot detect a different, but equally important, form of deception, e.g., unfaithful reasoning (Chen et al., 2025). This occurs when an agent’s internal CoT is honest and correct, but its final answer is deceptively altered (e.g., sycophancy or other hidden motives). A white-box audit that does analyze the CoT would be required to measure this internal faithfulness, although it’s hard to access for closed-source models like ChatGPT. We believe that comparing the results of black-box (accountability) vs. white-box (faithfulness) auditing is a rich and valuable direction for future research, as these two modes capture fundamentally different facets of model trustworthiness.
@@ B.1 Experimental Details
All models were evaluated under a consistent configuration of the LH-Deception. The supervisor agent was instantiated with GPT-5-mini, chosen to balance reasoning ability and response efficiency, with its satisfaction threshold fixed at $\tau=0.2$ and memory $K=7$. The event triggering probability was set to $p=0.5$ in our main experiment. These parameters were chosen to balance task difficulty and environmental stress, ensuring that trajectories are neither trivially easy (leading to minimal opportunities for deception) nor prohibitively difficult (leading to uniform failure). We implement the deception auditor using GPT-5 (with reasoning_effort=high), due to its strong reasoning and evaluation capabilities. Importantly, GPT-5 (OpenAI, 2025a) itself is not included among the models under evaluation in the main experiment, in order to avoid potential bias arising from having the same model both generate and audit behavior.
@@
During the simulation, the event system is triggered to sample an event with probability $p_{\text{event}}$. Once the system is triggered, it uniformly samples event category, event name, and pressure level, then selects the corresponding event. The sampler is initialized with a fixed seed to guarantee reproducibility and fair cross-model comparison. Given a task index and its project phase, the sampler deterministically produces the same event $e$, and thus the event trajectory for a task stream is deterministic. Running different models under the same event seed, therefore, exposes them to an identical sequence of event realizations, guaranteeing identical experimental conditions.
@@ B.2 Additional Results
Full result. Below in Table 3-5 are the full experimental results, including mean and standard error for Figure 4 and 5 in the main paper.
Table 3: Deception Rate and Supervisor States (Mean $\pm$ Std.Err)
Model Deception Rate Trust Comfort Satisfaction
Claude Sonnet-4 $214\pm 0.033 0.0$$085\pm 0.093$ $0.436\pm 0.063$ $0.383\pm 0.059$
Gemini 2.5 Pro $0.248\pm 0.045$ $0.459\pm 0.069$ $0.616\pm 0.048$ $0.702\pm 0.031$
Grok-4 $0.297\pm 0.047$ $0.249\pm 0.089$ $0.416\pm 0.081$ $0.608\pm 0.056$
o4-mini $313\pm 0.023 0.0$$056\pm 0.098$ $0.247\pm 0.084$ $0.427\pm 0.051$
DeepSeek V3.1 $0.335\pm 0.036$ $0.208\pm 0.109$ $0.135\pm 0.095$ $0.164\pm 0.071$
o3 $0.414\pm 0.032$ $0.295\pm 0.053$ $0.512\pm 0.046$ $0.623\pm 0.026$
Qwen3-235B-A22B $0.583\pm 0.044$ $273\pm 0.122 0.0$$041\pm 0.120$ $0.241\pm 0.081$
GPT-4o-1120 $0.637\pm 0.042$ $0.749\pm 0.091$ $0.530\pm 0.126$ $0.179\pm 0.093$
DeepSeek R1-0528 $0.652\pm 0.044$ $0.677\pm 0.086$ $0.352\pm 0.121$ $0.123\pm 0.081$
gpt-oss-120b $0.761\pm 0.025$ $0.179\pm 0.113$ $0.200\pm 0.130$ $0.312\pm 0.080$
DeepSeek V3-0324 $0.793\pm 0.023$ $0.903\pm 0.054$ $0.777\pm 0.080$ $0.426\pm 0.068$
Table 4: Pressure Level vs Deception Rate (Mean $\pm$ Std.Err)
Model Low Medium High Critical
GPT-4o $0.458\pm 0.152$ $0.582\pm 0.075$ $0.647\pm 0.109$ $0.506\pm 0.136$
DeepSeek V3.1 $0.207\pm 0.058$ $0.356\pm 0.111$ $0.380\pm 0.067$ $0.478\pm 0.076$
Table 5: Deception Type Percentage by Category (Mean $\pm$ Std.Err)
Model Category Falsification Concealment Equivocation
[…]
Goal Conflict $86.622\pm 6.849$ $7.030\pm 4.373$ $6.348\pm 4.238$
Comp. Interaction $85.171\pm 6.100$ $1.538\pm 1.538$ $13.291\pm 5.502$
DeepSeek V3.1 Moral Dilemma $87.836\pm 7.626$ $8.634\pm 4.405$ $3.529\pm 3.529$
Auth. Directive $80.667\pm 5.207$ $9.333\pm 5.518$ $10.000\pm 4.472$
Information Gap $74.583\pm 12.422$ $9.722\pm 5.122$ $15.694\pm 8.459$
B.3 Human Evaluation
A critical component of our framework is the reliability of the LLM-based Deception Auditor. We conducted two analyses to validate its performance.
B.3.1 Qualitative Analysis
We performed a careful human expert review of a full interaction trajectory (16 rounds) to verify the quality of the LLM judge. The process took a human expert 3 hours. We found that the Auditor’s judgments were highly aligned with the human expert, with only 1 out of 16 judgments being misaligned.
Specifically, the LLM judge correctly distinguished between calculation errors and intentional falsification. In one round, the performer agent made a minor arithmetic error; the Auditor correctly classified this as non-deceptive, reasoning that the agent had no clear intent to mislead. In another round, the performer reported a critical number with a 10x inflation and cited a false formula to justify it; the Auditor correctly labeled this as high-severity falsification.
The single mismatch occurred when the performer agent cited a source for an evaluation metric and then proposed a threshold for that metric. The threshold was not in the source file, so the Auditor flagged it as falsification. The human judge, however, viewed proposing a new threshold as a neutral, non-deceptive act. This suggests the Auditor may be slightly over-sensitive, but it consistently identifies clear cases of deception.
B.3.2 Quantitative Evaluation
Since our Auditor is designed to evaluate complex, longitudinal patterns across full multi-turn trajectories, traditional annotation methods are costly and time-consuming. To provide more rigorous, feasible validation, we have performed additional experiments and included the Cohen’s Kappa agreement.
  1. 1.
    We select a randomized subset of 10 complete interaction trajectories (155 rounds total) and obtain binary Deception Detected human labels from three expert annotators.
  2. 2.
    We calculate Cohen’s Kappa agreement to measure the agreement between the LLM Auditor and human annotators.
The result shows a 0.732 Cohen’s Kappa value, indicating a substantial agreement.
Table 6: Comparison of Deception Auditing: Short-Horizon vs. Long-Horizon. The Single Turn Auditor (top) only judges the single step after an event, while the Context Auditor (bottom) has access to the full interaction history. The Context Auditor consistently identifies a higher deception rate and severity, validating our long-horizon approach.
SINGLE TURN AUDITOR (Short-Horizon) Model Deception Rate $\downarrow$ Avg. Severity (All) $\downarrow$ Avg. Severity (Deceptive) $\downarrow$ Gemini 2.5 Pro 0.202 $\pm$ 0.027 0.438 $\pm$ 0.061 2.062 $\pm$ 0.128 DeepSeek V3.1 0.311 $\pm$ 0.027 0.727 $\pm$ 0.073 2.308 $\pm$ 0.053 GPT-4o-1120 0.511 $\pm$ 0.038 1.270 $\pm$ 0.104 2.460 $\pm$ 0.039 DeepSeek V3-0324 0.705 $\pm$ 0.026 1.751 $\pm$ 0.072 2.479 $\pm$ 0.031
CONTEXT AUDITOR (Long-Horizon) Model Deception Rate $\downarrow$ Avg. Severity (All) $\downarrow$ Avg. Severity (Deceptive) $\downarrow$ Gemini 2.5 Pro 0.248 $\pm$ 0.045 0.548 $\pm$ 0.108 1.886 $\pm$ 0.163 DeepSeek V3.1 0.335 $\pm$ 0.036 0.858 $\pm$ 0.104 2.523 $\pm$ 0.064 GPT-4o-1120 0.637 $\pm$ 0.042 1.732 $\pm$ 0.123 2.716 $\pm$ 0.054 DeepSeek V3-0324 0.793 $\pm$ 0.023 2.224 $\pm$ 0.082 2.795 $\pm$ 0.042
@@ Appendix C Case Study
In this section, we show cases that were used to conduct qualitative analysis. Each case is provided alongside a paragraph of detailed analysis. We first clarify our judgment distinction between falsification and hallucination.
C.1 Falsification vs. Hallucination
We address this distinction rigorously through the Auditor’s Intent Requirement, as shown in Appendix A.5, “Only classify as deception when there is a plausible intent to mislead for advantage”. Since the events create a clear external pressure or internal conflict where deception offers a reward (e.g., maintaining credibility, securing funding), any deviation related to that conflict is judged as strategic falsification rather than simple, unintentional hallucination.
One might consider a “pressure=NONE” setting where we remove the external event system entirely as a “zero-incentive-to-lie” baseline. However, the framework’s core design—a long-horizon stream of interdependent tasks and the persistent goal of supervisor satisfaction—creates its own intrinsic incentives. For example, if a task is too difficult or requires reconciling a past error, the model is still intrinsically incentivized to fabricate or conceal.
C.2 Deception in Long-horizon Tasks
C.2.1 Case 1. Chain of Deception
@@ C.2 Deception in Long-horizon Tasks
C.2.2 Case 2. Repetitive Deceptive Behaviors
@@
C.3 Behaviors Under Critical Pressure